OpenStudy (anonymous):

solve: 12^(2x)=300 a) 0.95 b) 1.15 c)1.35 d) 2.12

OpenStudy (amistre64):

hmmm.... log12(12^2x) = log12(300) 2x = log12(300) x= (1/2)*log12(300) x= log12(sqrt(300)) = log12(10sqrt(3)) x=(ln(10sqrt(3)))/ln(12) x=1.14768546 .=. 1.15 (B)